# On the semiclassical bounce with strong minimal assumptions

### v2

@@ Introduction
The mentioned conditions do not look completely impossible to satisfy but, at the present state of the art, it is not feasible to state that this kind of bounce is a realistic scenario to avoid singularity. Anyway, we can conclude that, in principle, the semiclassical effects in the radiation sector at the GUT scale may provide the singularity avoidance without additional ad hoc assumptions.
Another open question is the stability of the bounce model under discussion under the density and metric perturbations. This issue is typically complicated in all bounce models. The reason is that, in the vicinity of the bounce, the time derivative $\dot{H}$ is necessary positive and this leads to the violation of the Null Energy Condition (NEC). This feature may lead to instabilities in cosmic perturbations [38, 6] (see also [39] for an alternative discussion). In general, the analysis of the cosmological perturbations in a cosmological models is a necessary element of its development and this is especially true for models with a bounce [7]. Only the analysis of perturbations may show whether the given model is viable or possesses inconsistencies.
In the existing literature, there are strong indications of that the violation of NEC by quantum corrections may not lead to the inconsistencies [40] and that the same is true in the theories with scalar fields [41, 42]. Both arguments apply to our case. It is important to note that the perturbations should be analyzed not on the basis of a simplest non-covariant form of induced action (5), but using the covariant form (4), in the local representation. In this case, the theory always includes two auxiliary scalar fields [24, 25, 26] and, therefore, there are chances to arrive at the consistent model of bounce, including the perturbations free of pathologies, according to the criterion of [7]. We hope to come back to the detailed consideration of this issue and, as a first step, construct a new simplified formulation of the induced action with auxiliary fields, in a close future.
@@ Acknowledgments
Authors are very grateful to Patrick Peter for the interest to this work and very useful explanations concerning the role of cosmic perturbations in bounce models. W.C.S. wish to acknowledge CAPES for supporting his Ph.D. project. The work of I.Sh. is partially supported by Conselho Nacional de Desenvolvimento Científico e Tecnológico - CNPq under the grant 303635/2018-5.

### v3

@@ Introduction
The mentioned conditions do not look completely impossible to satisfy, but, at the present state of the art, it is not feasible to state that this kind of bounce is a realistic scenario to avoid singularity. Anyway, we can conclude that, in principle, the semiclassical effects in the radiation sector at the GUT scale may provide the singularity avoidance without additional ad hoc assumptions.
Another open question is the stability of the bounce model under discussion under the density and metric perturbations. This issue is typically complicated in all bounce models. The reason is that, in the vicinity of the bounce, the time derivative $\dot{H}$ is necessary positive, and this leads to the violation of the null energy condition (NEC). This feature may lead to instabilities in cosmic perturbations [38, 6] (see also [39] for an alternative discussion). The analysis of the cosmological perturbations in a cosmological model is a necessary element of its development, and this is especially true for models with a bounce [7]. Only the analysis of perturbations may show whether the given model is viable or possesses inconsistencies.
In general, additional restrictions on the bounce models come from the feature that the amplitude of the primordial power spectrum is usually proportional to the energy scale of the bounce. The power spectrum is constrained by CMB data, which can provide more stringent constraints on the bounce energy scale. This part was extensively discussed, e.g., in [40, 7, 41] and more recently in [42, 43]. The general situation is such that the mentioned constraints can be obtained only on the basis of the given cosmological model, as they are sensible to the structure of cosmic perturbations. [^3] The authors gratefully acknowledge discussions with Patrick Peter, Sandro Vitenti, and Alex Vikman on this subject. Thus, we leave the investigation of this issue to the next work with the analysis of perturbations.
In the existing literature, there are strong indications of that the violation of NEC by quantum corrections may not lead to the inconsistencies [44] and that the same is true in the theories with scalar fields [45, 46]. Both arguments apply to our case. It is important to note that the perturbations should be analyzed not on the basis of the simplest non-covariant form of induced action (5), but using the covariant form (4), in the local representation. In this case, the theory always includes two auxiliary scalar fields [24, 25, 26] and, therefore, there are chances to arrive at the consistent model of bounce, including the perturbations free of pathologies, according to the criterion of [7]. We hope to come back to the detailed consideration of this issue and, as a first step, construct a new simplified formulation of the induced action with auxiliary fields, in a close future.
@@ Acknowledgments
The authors are very grateful to Patrick Peter for his interest in this work and very useful explanations concerning the role of cosmic perturbations in bounce models. We also appreciate the help of Sandro Vitenti and Alex Vikman in explaining to us some details about CMB restrictions on the bounce models. W.C.S. thanks CAPES for supporting his Ph.D. project. The work of I.Sh. is partially supported by Conselho Nacional de Desenvolvimento Científico e Tecnológico - CNPq under Grant No. 303635/2018-5.